# A Spatial-statistical model to analyse historical rutting data

### v2

@@ 3.1 Likelihood: observed rutting given true rutting
3.2 Latent field: model for true rutting
We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$. Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$. We only consider models that have an interaction with AADT, and include a random yearly effect. It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables. This means that for pavement type, we assume that the rutting is linear for all pavement types (double traffic means double rutting), but that the different pavement types can have different rates of deterioration. From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions. To account for this, we include a random yearly effect in the model.
@@ 3.2.1 Non-spatial rutting model: based on explanatory variables and year effect
The model assumes that the level of expected rutting (mm/year) is influenced by an interaction effect of the annual average daily traffic (AADT) and some explanatory variables, including the level of rut depth from the previous year ($z_{\mathrm{d-1}}$), the lane width ($z_{\mathrm{w}}$) and pavement type. Technically, $z_{Ac}(s,t)$ is
$\begin{cases}\text{AADT (in ten thousands) if pavement at ($s,t$)is Ac,}\\ 0\ \text{if not Ac;}\\ \end{cases}$
similarly, for pavement types Agc and Sma, representing asphalt gravel concrete and stone mastic asphalt, respectively. The effect of lane width, say, on rutting is dependent on the traffic intensity. The random yearly effect is assumed to be distributed as $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},...,\gamma_{T}){\sim}\mathrm{Normal}(\boldsymbol{0}_{T},\ \tau_{\gamma}^{-1})$.
@@ 3.2.2 Models including spatial effects
There are several variables not included in our model (Equation (4)) that are known to influence rutting. Examples are variables related to weather, bearing capacity and proportion of heavy traffic load. Most of these variables are spatially dependent, and segments close in space are likely to be more similar than segments further a part. This can be accounted for and utilised by including spatial random effects. The spatial random effect is a Gaussian process (GP) along the road and can be interpreted as the part of the model not explained by the explanatory variables; lane width, rut depth from the previous year and pavement type that is spatially dependent. Further, it is reasonable that some of the spatial random effect is consistent over years (e.g., bearing capacity, side slope height, surface and base curvature and proportion of heavy traffic) while others vary between years (e.g., heavy rain events or more heavy traffic due to construction work). Therefore, we include both an annually varying spatial process $\xi_{t}(\boldsymbol{s})$ for year $t=1,...,T$, and a common spatial process $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component. The spatial rutting component is the spatial contribution to the total rutting that is not explained by the other variables e.g., variables in the non-spatial model in Equation (4). This contribution can either be positive or negative. The linear predictor for the model including spatial effects is described as
@@
The common spatial process $\omega(\boldsymbol{s})$ in Equations (5) is the time constant variable that can give rutting not explained by known variables (or unexplained rutting). It is a Gaussian random process $\left\{\omega(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\subseteq\mathbb{R}^{d}\right\}$ such that for any $n\geq 1$ and for each set of spatial locations $\left(s_{1},...,s_{n}\right)$ satisfies
@@ 4 Results and discussion
We fit the models in Section 3.2.3 to the data in Section 2.3. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation. The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance. Based on Model evaluation criteria (Table 5), and for easy comparison of the models, the spatial and non-spatial models with more explanatory variables are selected. The results for the remaining models are given in Appendix A.
@@ 4.1 Estimated spatial rutting component and yearly effect
We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, pavement type, and lane width. The spatial rutting component ()$\hat{\omega}(s)$) plotted in Figure 3(a), varies significantly along the road. Some locations exhibit up to $3.15$ mm/year extra rutting beyond what is accounted for by known factors, while others show rutting exceeding $2$ mm/year. Conversely, the minimum unexplained rutting is $-2.45$ mm/year, indicating better performance than expected.
Figure 3(b) presents estimates of the random yearly $\hat{\gamma}_{t}$. Rutting varies notably between years for spatial and non-spatial models, with minimum estimates in 2016 and maximum in 2018. A systematic shift between spatial and non-spatial models is evident, particularly in pavement type coefficients (see Table 3 in Section 4.2).
@@ 4.2 Estimated model parameters
The estimated parameters and 95% credible intervals (CI) for spatial (Model 1) and non-spatial (Model 4) are shown in Table 3. Spatial models exhibit notably higher estimates for pavement type coefficients compared with non-spatial models (see also Table A.2 in Appendix A). Moreover, models lacking the spatial rutting component display larger variance in the ‘noise’ term ($\hat{\sigma}{\varepsilon}$), suggesting that the spatial rutting component captures unobservable spatially varying factors such as weather and traffic.
@@ 4.2.1 Effects of the explanatory variables
Pavement types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3). Qian et al. (2019) and Mascarenhas et al. (2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on pavement types such as Agc and Ac. Notably, asphalt gravel concrete is predominantly situated approximately 30 kilometers from the Swedish border, where a sub-arctic climate, underdeveloped networks, and heavy traffic converge to adversely affect pavement performance Peel et al. (2007), Ketzler et al. (2021), Jenelius (2010), Hovi et al. (2018), Pinchasik et al. (2020). There is also greater uncertainty in asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), although considering spatial dependencies helps to explain more of the variability. Non-spatial models appear to explain variability between years rather than by pavement type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more crucial without accounting for spatial variability (Model 4, Table 3). The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is minimal and positive, suggesting negligible impact on expected rutting.
These findings indicate that changes in coefficients are significant when spatial dependencies of unmodeled factors (e.g., weather) are considered. Furthermore, the spatial model captures more uncertainty (smaller $\hat{\sigma}_{\varepsilon}$).
@@ 4.3 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
Refer to caption (c) $\hat{\eta}>2$ mm Refer to caption (d) $\hat{\eta}>3$ mm.
Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $*\hat{\eta}{{}^{*}}$, with each asphalt type highlighted. The effects are inferred for the year 2020. Graph (b): Expected rutting plotted against AADT and lane width for a short stretch. Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively.
@@ 4.3.1 Comparing expected rutting with observed data
In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning segments $1500-2700$ (or $30-54$ km from Stjørdal), characterized by weaker pavement type Agc. This section exhibits spikes of accelerated rutting, particularly evident with increasing traffic. Possible factors contributing to this variability include changes in asphalt type, traffic composition, suboptimal road design, or inadequate pavement maintenance.
@@ 4.4 Investigating an area with high rutting rates
The significant spike in Figure (4) (a), indicating high rutting rates, prompts further investigation. We examine road segments spanning 2100 to 2700 meters surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien. This stretch, depicted in Figure 4 (a), exhibits notably high rutting, peaking at $4.67$ mm, with an estimated lifespan of only $5$ years (in a low-traffic area requiring maintenance at 25 mm rut depth). Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches (see Figure 5), heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4). Meråker’s industrial and agricultural activities result in increased traffic from heavy vehicles, including farm equipment such as tractors (industrial, wheel drive, row crop).
(a) Short section between EV14 S6D1 m3841-m4285 F1 with high rutting.
(b) Location EV14 S6D1 m4068 F1 in year 2020
(c) Location EV14 S6D1 m4059 F1 in year 2021
(d) Location EV14 S6D1 m4065 F1 in year 2022
Figure 5: Graph (a): Location with high rutting rates (road segments $2100-2700$ towards Sweden). Graphs (b), (c) and (d) show the location, in years 2020-2022 (see Statens vegvesen or url: https://vegkart.atlas.vegvesen.no/#kartlag:nib/@335074,7038866,15/vegsystemreferanse:335035.067:7038844.352
@@ 4.4.1 Exploring the spatial model with various values of the explanatory variables
To further investigate high rutting values in a road section, we analyze the impact of explanatory variables, yearly variation, and spatial rutting component ($\hat{\eta}$). We explore spatial Model 1 using different values for variables to estimate rutting. First, we use actual measurements like asphalt type, lane width, and rut depth ($\hat{\eta}_{\mathrm{True}}$). Second, we consider alternative asphalt types - asphalt concrete (Ac) or stone mastic asphalt (Sma) ($\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$). Third, we examine extreme values such as the strongest pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$). Conversely, accelerated rutting scenarios involve deeper rut depth (25 mm), narrower lanes (2.75 meters), and asphalt gravel concrete (Agc) ($\hat{\eta}_{\mathrm{Worse}}$) from the E14 dataset. These scenarios are depicted in Figure 6 (a).
@@
Using the best covariate values ( ), rutting is projected to be at most 4.30 mm/year, with maintenance required approximately 6 years after construction or maintenance. Additionally, using the stronger asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor. Asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) show similar rutting estimates due to similar binder concentrations.
(a) Expected rutting at road segments $2100-2500$ (metres) towards Sweden.
@@
Figure 6: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road segments $2100-2700$, along with the ’best’ or ’worst’ explanatory values. $\hat{\eta}_{\mathrm{Agc}}$ represents pavement type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$. Graph (b) illustrates expected rutting inferred from $\hat{\omega}(s)$, $\hat{\gamma}$, and sequentially added explanatory variables. Graph (c) shows 2021 bearing capacity in tons; values $\geq$ 16 indicate good parameters. Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14; values $<$ 12 indicate good curvature index.
We examine the impact of pavement type on rutting (Figure 6, graph (b)). Changing from asphalt type Agc (asphalt gravel concrete) to the stronger Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with pavement type. Initially, pavement type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Subsequently, the previous year’s rut depth is incorporated, represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Similarly, the model with added lane width ${\hat{\beta}_{w}z_{\mathrm{w(s,t)}}}$ is denoted as $\hat{\tilde{\eta}}_{\mathrm{w}}$. Results from using stone mastic asphalt (Sma) are depicted by solid lines, and asphalt gravel concrete (Agc) results are shown with dotted lines.
The findings affirm that pavement type significantly influences rutting. Implementing the stronger binder type Sma instead of the actual pavement type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually. Over a projected lifetime of 6 years, this equates to an additional year of serviceability.

### v3

@@ 3.1 Likelihood: observed rutting given modelled rutting
3.2 Latent field: modelled rutting
We model the observed rutting as Gaussian with expected value $\eta_{t}(s_{i})$ and variance $\sigma^{2}_{\varepsilon}$. Further, $r_{i,t}$ is conditionally independent, given the latent field $\eta_{t}(s_{i})$. We only consider models that have an interaction with AADT, and include a random yearly effect. It is known that more traffic gives more rutting, so we include traffic (AADT) in the model as an interaction with each of the explanatory variables. This means that for asphalt type, we assume that the rutting is linear for all asphalt types (double traffic means double rutting), but that the different asphalt types can have different rates of deterioration. From the explanatory analysis (Figure 2), we found that the rutting seems to be on different levels for different years, and this might be explained by different weather conditions. To account for this, we include a random yearly effect in the model.
@@ 3.2.1 Non-spatial rutting model: based on explanatory variables and year effect
The model assumes that the level of expected rutting (mm/year) is influenced by an interaction effect of the annual average daily traffic (AADT) and some explanatory variables, including the level of rut depth from the previous year ($z_{\mathrm{d-1}}$), the lane width ($z_{\mathrm{w}}$) and asphalt type. This assumption stems from our prior understanding: when there’s no traffic, there’s no rutting (or rut depth). This insight is grounded in research, such as the work of Hao et al. (2009), which identified increased loading time (traffic loads) as the primary factor accelerating rutting in asphalt pavements.
Let $z_{\mathrm{AADT}(s,t)}$ denote AADT for location $s$ in year $t$, and $I_{\mathrm{Ac}(s,t)}$ be an indicator variable that has a value of 1 if location $s$ in year $t$ has asphalt type Ac and 00 otherwise. Then, $z_{\mathrm{Ac}(s,t)}=I_{\mathrm{Ac}(s,t)}\cdot z_{\mathrm{AADT}(s,t)}$ and, hence, equals AADT if the asphalt type is Ac (asphalt concrete) and 00 otherwise. Similarly, $z_{\mathrm{Agc}(s,t)}$ and $z_{\mathrm{Sma}(s,t)}$ equal the AADT for location $s$ in year $t$ if the asphalt type is Agc (asphalt gravel concrete) and Sma (stone mastic asphalt), respectively, and 00 otherwise. Further, $z_{\mathrm{d-1}(s,t)}=d(s,t)\cdot z_{\mathrm{AADT}(s,t)}$ and $z_{\mathrm{w(s,t)}}=\mathrm{w}{(s,t)}\cdot z_{\mathrm{AADT}(s,t)}$.
The random yearly effect is assumed to be distributed as $\boldsymbol{\gamma}=(\gamma_{1},\gamma_{2},...,\gamma_{T}){\sim}\mathrm{Normal}(\boldsymbol{0}_{T},\ \tau_{\gamma}^{-1})$.
@@ 3.2.2 Models including spatial effects
Our model (Equation (3)) does not include several variables known to influence rutting, such as weather, bearing capacity, and proportion of heavy traffic load. These variables are spatially dependent, with segments closer in space likely to be more similar than those further apart. To account for this spatial dependency, we incorporate a Gaussian process (GP) as a spatial random effect along the road. GPs are powerful tools for modeling spatial dependencies, capturing variability due to both known factors (e.g., road characteristics, traffic patterns, and other explanatory variables) and unknown factors not explicitly accounted for in the model. The unknown factors could include unmeasured road characteristics, environmental variables like temperature, or other spatially dependent factors influencing rutting. Our model aims to capture both consistent spatial effects that remain stable over time and spatial effects that vary between years.
Consistent spatial effects, such as road characteristics (e.g., bearing capacity, side slope height, surface and base curvature, and proportion of heavy traffic), are likely to influence rutting consistently over time. Additionally, there are spatial effects that vary between years, such as heavy rain events, temperature variations, or fluctuations in heavy traffic due to construction work. To account for these variations, we include both an annually varying spatial process, $\xi_{t}(\boldsymbol{s})$, for each year $t=1,...,T$, and a common spatial process, $\omega(\boldsymbol{s})$, referred to as the spatial rutting component.
The annually varying spatial process, $\xi_{t}(\boldsymbol{s})$, captures the spatial variability in rutting that changes from year to year, accounting for spatial variations specific to each year, such as the impact of specific weather events or construction activities.
The common spatial process, $\omega(\boldsymbol{s})$, which we refer to as the spatial rutting component, captures the spatial variability in rutting that remains consistent across all years. This includes spatial effects not explicitly accounted for by other variables in the model, such as road characteristics and other spatially dependent factors. This contribution can be either positive or negative. The linear predictor for the model including spatial effects is described as
@@
The common spatial process $\omega(\boldsymbol{s})$ in Equations (4) is the time constant variable. It is a Gaussian random process $\left\{\omega(\boldsymbol{s}):\boldsymbol{s}\in\mathcal{D}\subseteq\mathbb{R}^{d}\right\}$ such that for any $n\geq 1$ and for each set of spatial locations $\left(s_{1},...,s_{n}\right)$ satisfies
@@ 4 Results and discussion
We fit the models in Section 3.2.3 to the data in Section LABEL:sec:variables. For interpretability, the explanatory variables are standardized by subtracting the mean and dividing by the standard deviation. The results from the spatial Model 1 and the non-spatial version Model 4 are used to demonstrate the models’ performance. Based on Model evaluation criteria (Table 5), and for easy comparison of the models, the spatial and non-spatial models with more explanatory variables are selected. The results for the remaining models are given in Appendix A.
@@ 4.1 Estimated spatial rutting component and yearly effect
We analyze estimates from Model 1, identifying areas with high rutting rates not explained by AADT, asphalt type, and lane width. The spatial rutting component ()$\hat{\omega}(s)$) plotted in Figure 3(a), varies significantly along the road. Some locations exhibit up to $3.15$ mm/year extra rutting beyond what is accounted for by known factors, while others show rutting exceeding $2$ mm/year. Conversely, the minimum unexplained rutting is $-2.45$ mm/year, indicating better performance than expected.
Figure 3(b) presents estimates of the random yearly $\hat{\gamma}_{t}$. Rutting varies notably between years for spatial and non-spatial models, with minimum estimates in 2016 and maximum in 2018. A systematic shift between spatial and non-spatial models is evident, particularly in asphalt type coefficients (see Table 3 in Section 4.2).
@@ 4.2 Estimated model parameters
The estimated parameters and 95% credible intervals (CI) for spatial (Model 1) and non-spatial (Model 4) are shown in Table 3. Spatial models exhibit notably higher estimates for asphalt type coefficients compared with non-spatial models (see also Table A.2 in Appendix A). Moreover, models lacking the spatial rutting component display larger variance in the ‘noise’ term ($\hat{\sigma}{\varepsilon}$), suggesting that the spatial rutting component captures unobservable spatially varying factors such as weather and traffic.
@@ 4.2.1 Effects of the explanatory variables
Asphalt types, particularly those with weaker binder concentrations like asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) and asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$), experience more rutting (Table 3). Qian et al. (2019) and Mascarenhas et al. (2020) shed light on how temperature and heavy traffic ($>$ 3.5 tonnes) impact rutting resistance on asphalt types such as Agc and Ac. Notably, asphalt gravel concrete is predominantly situated approximately 30 kilometers from the Swedish border, where a sub-arctic climate, underdeveloped networks, and heavy traffic converge to adversely affect pavement performance Peel et al. (2007), Ketzler et al. (2021), Jenelius (2010), Hovi et al. (2018), Pinchasik et al. (2020). There is also greater uncertainty in asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), although considering spatial dependencies helps to explain more of the variability. Non-spatial models appear to explain variability between years rather than by asphalt type, suggesting a nuanced relationship between these factors in understanding rutting susceptibility.
The rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes increasingly important when spatial variability is not considered (Model 4, Table 3). As a genuine spatial-temporal variable, including the previous year’s rut depth in the model modifies the spatial effect $\hat{\sigma}_{\omega}(s)$, specifically reducing it, which in turn enhances the precision of the estimates by decreasing uncertainty, $\hat{\sigma}_{\varepsilon}$. This improvement is evident in Models 1 and 3 (see Table 4 in Section A). Incorporating additional relevant spatial-temporal variables could further refine the results, providing a better explanation for the observed uncertainty.
The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is minimal and positive, suggesting negligible impact on expected rutting.
@@ 4.3 Estimated rutting from the explanatory variables, random yearly effect and spatial rutting component
(c) $\hat{\eta}>2$ mm
(d) $\hat{\eta}>3$ mm.
Figure 4: Graph (a): Expected rutting from: i) $\hat{\eta}$ and ii) from $\hat{\eta}{{}^{*}}$, with each asphalt type highlighted. The effects are inferred for the year 2020. Graph (b): Expected rutting plotted against AADT and lane width for a short stretch. Graphs (c) and (d) give the locations where rutting exceeds $2$ mm/year and $3$ mm/year, respectively
@@ 4.3.1 Comparing expected rutting with observed data
In Figure 4, graph (b), we plot the expected rutting $\hat{\eta}$, lane width, and AADT (per ten thousand) for a short road section spanning road stretch 42000 to 54000 metres (or $42-54$ km from Stjørdal), characterized by weaker asphalt type Agc. This section exhibits spikes of accelerated rutting, particularly evident with increasing traffic. Possible factors contributing to this variability include changes in asphalt type, traffic composition, suboptimal road design, or inadequate pavement maintenance.
@@ 4.4 Investigating an area with high rutting rates
The significant spike in Figure (4) (a), indicating high rutting rates, prompts further investigation. We examine the road stretch spanning 42000 to 54000 metres surrounding this spike, located in the municipality of Meråker, just 20 kilometers west of the Swedish border at Storlien. This stretch, depicted in Figure 4 (a), exhibits notably high rutting, peaking at $4.67$ mm, with an estimated lifespan of only $5$ years (in a low-traffic area requiring maintenance at 25 mm rut depth). Possible contributors to this accelerated rutting include clogged drainage, flooding, absence of side ditches, heavy traffic, or road designs featuring railguards installed close to the lane edge, which encourages traffic to remain within the same wheel tracks (refer to Figure A.4 in Appendix A.4). Meråker’s industrial and agricultural activities result in increased traffic from heavy vehicles, including farm equipment such as tractors (industrial, wheel drive, row crop).
@@ 4.4.1 Exploring the spatial model with various values of the explanatory variables
To further investigate high rutting values in a road section, we analyze the impact of explanatory variables, yearly variation, and spatial rutting component ($\hat{\eta}$). We explore spatial Model 1 using different values for variables to estimate rutting. First, we use actual measurements like asphalt type, lane width, and rut depth ($\hat{\eta}_{\mathrm{True}}$). Second, we consider alternative asphalt types - asphalt concrete (Ac) or stone mastic asphalt (Sma) ($\hat{\eta}_{\mathrm{Ac}}$ or $\hat{\eta}_{\mathrm{Sma}}$). Third, we examine extreme values such as the more wear- and deformation-resistant pavement (Sma), minimal rut depth (0.01 mm), or wider lanes (6 meters) ($\hat{\eta}_{\mathrm{Best}}$). Conversely, accelerated rutting scenarios involve deeper rut depth (25 mm), narrower lanes (2.75 meters), and asphalt gravel concrete (Agc) ($\hat{\eta}_{\mathrm{Worse}}$) from the E14 dataset. These scenarios are depicted in Figure 5 (a).
@@
Using the best covariate values ( ), rutting is projected to be at most 4.30 mm/year, with maintenance required approximately 6 years after construction or maintenance. Additionally, using the asphalt binder Sma yields similar rutting estimates but generally lower than those using the best values, indicating asphalt type as a primary factor. Asphalt gravel concrete (Agc,  ) and asphalt concrete (Ac,  ) show similar rutting estimates due to similar binder concentrations.
(a) Expected rutting at road segments $42000-54000$ (metres) towards Sweden.
@@
Figure 5: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road stretch 42000 to 54000 metres, along with the ‘best’ or ‘worst’ explanatory values. $\hat{\eta}_{\mathrm{Agc}}$ represents asphalt type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$. Graph (b) illustrates expected rutting inferred from $\hat{\omega}(s)$, $\hat{\gamma}$, and sequentially added explanatory variables. Graph (c) shows 2021 bearing capacity in tons; values $\geq$ 16 indicate good parameters. Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14, where values $<$ 12 indicate good curvature index. Maximum y-value plotted is 50 (values $>50$ are denoted by the red ‘x’ symbol).
We examine the impact of asphalt type on rutting (Figure 5, graph (b)). Changing from asphalt type Agc (asphalt gravel concrete) to the more wear- and deformation-resistant Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with asphalt type. Initially, asphalt type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Subsequently, the previous year’s rut depth is incorporated, represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Similarly, the model with added lane width ${\hat{\beta}_{w}z_{\mathrm{w(s,t)}}}$ is denoted as $\hat{\tilde{\eta}}_{\mathrm{w}}$. Results from using stone mastic asphalt (Sma) are depicted by solid lines, and asphalt gravel concrete (Agc) results are shown with dotted lines.
The findings affirm that asphalt type significantly influences rutting. Implementing the binder type Sma instead of the actual asphalt type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually. Over a projected lifetime of 6 years, this equates to an additional year of serviceability.